The relationship between the displacement S of a body and
time t is given by the equation: S = A − Bt + Ct2. S is expressed
in meter and t in seconds. Numerical values of the constants are
A = 6,B = 3 and C = 1 with appropriate SI units. Mass of the
body is 4 kg. If the motion begins at t = 0, calculate the power
delivered to the body at t = 4 seconds. Express your answer in
watt.

Grade:10

1 Answers

MAINAK KHAN
6 years ago
POWER= FORCE* VELOCITY
POWER= MASS*Acceleration*VELOCITY
S=displacement
v=velocity
a= acceleration
ds/dt=velocity
d(A-Bt+Ct2)/dt= -B+2Ct
v=-B+2Ct
v=-3+2(1)4                                         (since power at t=4sec)
v=5m/s
dv/dt=acceleration
d(-B+2Ct)/dt=+2C
a=2C
a=2(1)=2m/s2
POWER=4(2)5=40watt